# GenOM: Ontology Matching with Description Generation and Large Language Model

### v1

@@ 4.4.1 Impact of Definition Enrichment on Local Ranking and Candidate Retrieval
We study the impact of the generated concept definition on each of the two stages: LLM-based equivalence judgement and candidate generation. For LLM-based equivalence judgement, we compare using only the concept label, and using both the label and the generated definition. The local ranking result is shown in Table 6. In particular, both MRR and Hit@1 show noticeable gains in all tasks except for SNOMED-FMA (body), where performance remains comparable.
For the candidate generation stage, we additionally report Hit@5 and Hit@10, as these metrics are essential for determining the appropriate top-$k$ value—that is, how many candidate concepts should be passed to the LLM for equivalence judgement. As shown in Table 7, incorporating definition information led to improvements across all three metrics (Hit@1, Hit@5, Hit@10) in all tasks, except for a slight drop in Hit@10 on the SNOMED-NCIT (pharm) task. These results suggest that enriched definitions help retrieve a greater number of correct candidates, thereby increasing the likelihood of including the true match within the top-$k$ shortlist.
Task MRR Hit@1
OMIM_ORDO (with) 0.880 0.831
OMIM_ORDO (without) 0.659 0.543
NCIT_DOID (with) 0.889 0.820
NCIT_DOID (without) 0.858 0.782
SNOMED_NCIT_pharm (with) 0.853 0.770
SNOMED_NCIT_pharm (without) 0.777 0.688
SNOMED_NCIT_neoplas (with) 0.863 0.777
SNOMED_NCIT_neoplas (without) 0.824 0.732
SNOMED_FMA_body (with) 0.754 0.631
SNOMED_FMA_body (without) 0.764 0.633
@@
Table 6: Impact of concept definitions on LLM-based Local Ranking (Qwen-2.5-7B)
Task Hit@1 Hit@5 Hit@10
OMIM_ORDO (with) 0.737 0.879 0.909
OMIM_ORDO (without) 0.724 0.875 0.907
NCIT_DOID (with) 0.886 0.974 0.985
NCIT_DOID (without) 0.786 0.940 0.962
SNOMED_NCIT_pharm (with) 0.747 0.942 0.966
SNOMED_NCIT_pharm (without) 0.712 0.911 0.946
SNOMED_NCIT_neoplas (with) 0.722 0.896 0.932
SNOMED_NCIT_neoplas (without) 0.718 0.896 0.938
SNOMED_FMA_body (with) 0.695 0.912 0.947
SNOMED_FMA_body (without) 0.642 0.883 0.930
@@
Table 7: Impact of concept definitions on embedding-based candidate retrieval (text-embedding-3-small)
4.4.2 Effectiveness Compared to Original Exact Matching Methods
We compare the results of GenOM with the results of the stand-alone exact matching system that GenOM adopts in the final stage (either BERTMapLt or LogMapLt). The BERTMapLt and LogMapLt results reported in this section are reproduced within the scope of this work, and may therefore differ slightly from the results presented earlier. As shown in Table 8, GenOM consistently outperforms the original exact matchers across all evaluated tasks.
GenOM (BERTMapLt) achieved an average F1 score of 0.771 across the five benchmark tasks, outperforming the original BERTMapLt model, which obtained an average of 0.744. A similar improvement was observed with LogMapLt: while the standalone LogMapLt achieved an average F1 score of only 0.631, the GenOM-enhanced version reached 0.716. These results suggest that while exact matching provides a solid foundation for identifying high-confidence correspondences, it remains limited in capturing more nuanced semantic equivalence. By integrating LLM-based reasoning and enriched conceptual representations, GenOM is able to significantly enhance both coverage and accuracy over the base exact matching techniques. This is reflected in a notable increase in recall: GenOM achieves, on average, an 8% improvement in recall over BERTMapLt, and an even more substantial 24% increase when compared to LogMapLt.
@@ 4.4.2 Effectiveness Compared to Original Exact Matching Methods
SNOMED-NCIT-neoplas GenOM(BERTMapLt) 0.795 0.764 0.779
@@
SNOMED-NCIT-pharm GenOM(BERTMapLt) 0.989 0.600 0.747
@@
SNOMED-FMA-body GenOM(BERTMapLt) 0.944 0.678 0.789
@@
NCIT-DOID GenOM(BERTMapLt) 0.912 0.846 0.878
@@
OMIM-ORDO GenOM(BERTMapLt) 0.803 0.565 0.664
BERTMapLt 0.834 0.497 0.623
GenOM(LogMapLt ) 0.839 0.394 0.537
LogMapLt 0.937 0.215 0.350
Table 8: The results of GenOM and the exact matching systems BERTMapLt and LogMapLt. The results of BERTMapLt and LogMapLt here are reproduced, as a component as of the GenOM framework, and have a small difference as those in Table 5.
4.4.3 Effectiveness of Few-Shot Prompting
This experiment also investigates the effect of few-shot prompting on the LLM-based equivalence judgement stage. The results are shown inTable 9, where few-shot prompting is set to 2 examples, all results are reported prior to the integration of exact matching, and the threshold for cosine similarity was kept consistent with the earlier setting at 0.97. Aside from the inclusion of few-shot examples, all other settings are identical. The evaluation was conducted without incorporating results from the exact matching module, as the impact of the few-shot strategy tends to be diminished once exact matching is applied. The results indicate that incorporating two-shot examples consistently improves performance across most tasks. Except for the NCIT-DOID task, where performance remains unchanged, all other tasks exhibit notable gains in F1 score. This demonstrates that few-shot prompting can effectively guide the LLM towards more accurate classification, especially in borderline cases where single-instance reasoning may be insufficient.
Task Prompt Strategy P R F1
SNOMED-NCIT-neoplas LLM (Qwen2.5 7B) 0.820 0.308 0.448
Few-shot 0.815 0.346 0.486
SNOMED-NCIT-pharm LLM (Qwen2.5 7B) 0.94 0.183 0.306
Few-shot 0.919 0.21 0.342
SNOMED-FMA-body LLM (Qwen2.5 7B) 0.789 0.016 0.113
Few-shot 0.846 0.166 0.278
NCIT-DOID LLM (Qwen2.5 7B) 0.942 0.474 0.631
Few-shot 0.935 0.476 0.632
OMIM-ORDO LLM (Qwen2.5 7B) 0.807 0.257 0.390
Few-shot 0.795 0.271 0.405
Table 9: The results of the LLM-based equivalence judgement stage of GenOM, with and without few-shot prompting
@@ 5 Conclusion, Discussion and Future Work
This paper presents GenOM, a general-purpose framework for ontology alignment that integrates concept semantic enrichment with LLM-based textual definition generation, embedding-based candidate retrieval, LLM prompting-based equivalence judgement, and exact matching in a modular design. The approach demonstrates strong performance across five biomedical ontology alignment tasks of OAEI Bio-ML, outperforming many baselines, and the effectiveness of its important components has been verified via extensive ablation studies. In particular, the framework shows its ability to generalise across datasets while maintaining alignment accuracy, without relying on handcrafted features or extensive task-specific engineering.
@@
    The definition of equivalence which can vary subtly across tasks: concept pairs deemed equivalent in one alignment task may not be considered so in another, leading to inconsistencies in judgement. Accordingly, the optimal similarity threshold becomes task-dependent. For example, a cosine similarity above 0.80 indicates equivalence in some tasks, but some other tasks may require a threshold of 0.95 to ensure equivalence.
  3. 3.
@@
    The alignment performance of LLMs is highly sensitive to the prompt. In evaluation, we observed that vague prompts such as simply asking the model to “determine whether two concepts are equivalent” often fails to elicit correct predictions; in many cases, the LLM almost never produces a “YES” output. This highlights the importance of prompt specificity in steering LLM behaviour and underscores a practical challenge in applying LLMs to alignment in a generalisable way.
@@
For the future work, one key direction is to expand the scope of GenOM to include additional alignment types beyond equivalence, such as subsumption. Another key direction involves addressing the variability in how equivalence is defined across different ontologies and tasks. In many alignment scenarios, the threshold for considering two concepts equivalent may depend on contextual or domain-specific nuances, which are difficult to capture using a fixed similarity score or binary decision. To tackle this, future research will explore task-adaptive alignment criteria, including dynamic threshold selection and prompt-based calibration techniques that allow the LLM to assess the strength or type of correspondence more flexibly. Additionally, incorporating finer-grained semantic similarity measures and confidence estimation strategies could help better reflect the spectrum of equivalence relations observed in practice.

### v2

@@ 7.1 Impact of Definition Enrichment on Local Ranking and Candidate Retrieval
We investigated the impact of definition-based semantic enrichment on two stages of the pipeline: LLM-based equivalence judgement and embedding-based candidate generation. For LLM-based equivalence judgement, we compared a label-only setting with a label-plus-definition setting. The local ranking result was shown in Figure 4. In particular, both MRR and Hit@1 showed clear improvements in all tasks except for SNOMED-FMA (body), where performance remains comparable.
OMIM–ORDONCIT–DOIDSNOMED–NCIT (pharm)SNOMED–NCIT (neoplas)SNOMED–FMA (body)0.60.60.80.811MRRWith definitionsWithout definitions
(a) MRR
OMIM–ORDONCIT–DOIDSNOMED–NCIT (pharm)SNOMED–NCIT (neoplas)SNOMED–FMA (body)0.60.60.80.811Hit@1
(b) Hit@1
@@
Figure 4: Impact of concept definitions on LLM-based local ranking (Qwen32B).
For the embedding-based candidate generation stage, we additionally reported Hit@5 and Hit@10, as these metrics are essential for determining the appropriate top-$k$ value—that is, how many candidate concepts should be passed to the LLM for equivalence judgement. As shown in Figure 5, incorporating definition information led to improvements across all three metrics (Hit@1, Hit@5, Hit@10) in all tasks, except for a slight drop in Hit@10 on the SNOMED-NCIT (pharm) task. These results suggested that enriched definitions help retrieve a greater number of correct candidates, thereby increasing the likelihood of including the true match within the top-$k$ shortlist.
OMIM–ORDONCIT–DOIDSNOMED–NCIT (pharm)SNOMED–NCIT (neoplas)SNOMED–FMA (body)0.60.60.70.70.80.80.90.911Hit@1With definitionsWithout definitions
(a) Hit@1
OMIM–ORDONCIT–DOIDSNOMED–NCIT (pharm)SNOMED–NCIT (neoplas)SNOMED–FMA (body)0.60.60.70.70.80.80.90.911Hit@5
(b) Hit@5
OMIM–ORDONCIT–DOIDSNOMED–NCIT (pharm)SNOMED–NCIT (neoplas)SNOMED–FMA (body)0.60.60.70.70.80.80.90.911Hit@10
(c) Hit@10
@@
Figure 5: Impact of concept definitions on embedding-based candidate retrieval (text-embedding-3-small).
7.2 Effectiveness Compared to Original Exact Matching Methods
@@
SNOMED-NCIT-neoplas GenOM(BERTMapLt) 0.747 0.804 0.774
@@
SNOMED-NCIT-pharm GenOM(BERTMapLt) 0.957 0.622 0.754
@@
SNOMED-FMA-body GenOM(BERTMapLt) 0.902 0.674 0.771
@@
OMIM-ORDO GenOM(BERTMapLt) 0.834 0.643 0.726
BERTMapLt 0.834 0.497 0.623
GenOM(LogMapLt) 0.886 0.544 0.674
LogMapLt 0.937 0.215 0.350
NCIT-DOID GenOM(BERTMapLt) 0.866 0.937 0.901
@@
Table 8: The results of GenOM with Qwen32B using the exact matching systems BERTMapLt and LogMapLt, as well as the results of BERTMapLt and LogMapLt alone.
We compared the results of GenOM (Qwen32B) with the results of the stand-alone exact matching system that GenOM (Qwen32B) adopted in the final stage (either BERTMapLt or LogMapLt). The BERTMapLt and LogMapLt results reported in this section were reproduced within the scope of this work and therefore might differ slightly from the results presented in Table 5.
As shown in Table 8, GenOM (Qwen32B) surpassed the strongest exact-matching systems on most tasks, with the sole exception of the SNOMED–FMA (Body) alignment. In our earlier workshop version of this work [13], which focused on 7B-level LLMs, we were able to outperform both BERTMapLt and LogMapLt across all tasks by imposing a relatively high cosine-similarity threshold. However, as discussed previously, larger LLMs exhibited much lower sensitivity to the cosine-similarity threshold, with Qwen-32B showing only minor fluctuations in F1 scores between thresholds of 0.6 and 0.9. Under such circumstances, enforcing extremely strict thresholds (e.g. 0.97), while effective for smaller models, is less suitable for larger ones, making a more balanced configuration preferable.
Even with this trade-off, GenOM (Qwen32B) consistently improved upon exact matching across the Bio-ML tasks. On average, GenOM (Qwen32B) achieved a 5.5% relative F1 improvement over BERTMapLt and a more substantial 15.7% improvement over LogMapLt. These gains demonstrated that the proposed framework remained effective even without relying on aggressive threshold filtering, and that the enhanced reasoning capabilities of larger LLMs could be translated into measurable performance improvements over strong symbolic baselines.
@@ 8 Conclusion, Discussion and Future Work
This paper presents GenOM, a general-purpose framework for ontology alignment that integrated concept semantic enrichment with LLM-based textual definition generation, embedding-based candidate retrieval, LLM prompting-based equivalence judgement, and exact matching in a modular design. Evaluation results have shown the approach demonstrates strong performance across five biomedical ontology alignment tasks of OAEI Bio-ML, outperforming many baselines, and the effectiveness of its important components has been verified via ablation studies. In particular, the framework shows its ability to generalise across datasets while maintaining alignment accuracy, without relying on handcrafted features or extensive task-specific engineering.
@@
    The alignment performance of LLMs is highly sensitive to the prompt. In the evaluation, we observed that vague prompts such as simply asking the model to “determine whether two concepts are equivalent” often fails to elicit correct predictions; in many cases, the LLM almost never produces‘’YES’ output. This highlights the importance of prompt specificity in steering LLM behaviour and underscores a practical challenge in applying LLMs to alignment in a generalisable way.
@@
For future work, an important enhancement would be to expand the scope of GenOM to include additional alignment types beyond equivalence, such as subsumption. Another key direction would involves addressing the variability in how equivalence is defined across different ontologies and tasks. In many alignment scenarios, the threshold for considering two concepts equivalent may depend on contextual or domain-specific nuances, which are difficult to capture using a fixed similarity score or binary decision. To tackle this, task-adaptive alignment criteria would be useful to explore, including dynamic threshold selection and prompt-based calibration techniques that allow the LLM to assess the strength or type of correspondence more flexibly. Additionally, incorporating finer-grained semantic similarity measures and confidence estimation strategies could help better reflect the spectrum of equivalence relations observed in practice.
Appendix A
Table 9 reports the per-task optimal performance that GenOM achieves when the decision thresholds are tuned separately for each ontology matching task. As expected, task-specific tuning leads to further improvements over the unified threshold setting reported in the main paper, reflecting the heterogeneous similarity distributions across different Bio-ML tracks. The results are not used for comparison in the study and do not replace the unified-threshold setting reported in the main text. They are provided as auxiliary evidence about the potential performance of GenOM(Qwen32B) under task-specific calibration when a set of uniformly sampled mappings are annotated for validation and hyper parameter searching.
Table 9: Per-task optimal performance of GenOM(Qwen32B) under task-specific thresholds (Precision, Recall, and F1-score).
Bio-ML Track Precision Recall F1-score
SNOMED-NCIT (Neoplas) 0.812 0.765 0.788
SNOMED-NCIT (Pharm) 0.915 0.653 0.762
SNOMED-FMA (Body) 0.921 0.699 0.795
OMIM-ORDO 0.824 0.705 0.760
NCIT-DOID 0.873 0.938 0.904
Appendix B
B.1 Qualitative Sanity Check via Manual Inspection
As a qualitative sanity check, we conducted a follow-up manual inspection of the LLM-based evaluation outputs. Due to practical resource constraints, this inspection was limited to two datasets, Pharm-SNOMED and Body-FMA, covering a total of 200 biomedical concepts. The inspection was carried out by the authors using external resources, including publicly available biomedical references and search engines, and was further supplemented by consultation with general-purpose LLMs to assist in understanding unfamiliar medical concepts. Although the authors are not domain experts and did not examine all concepts evaluated by the LLM-based scheme, we observed a high degree of consistency between the manual assessments and the score produced by the automated evaluation within the inspected subset. This analysis was not intended as a formal validation, but rather as an auxiliary plausibility check to assess whether the proposed evaluation method produced rankings that were manifestly counter-intuitive.
B.2 Prompt Template and Analysis
During experimentation, we observed that the LLM occasionally interchanged the numerical values assigned to discriminative and correctness in its JSON output. This issue reflects a known limitation of LLMs in adhering to strict output schemas, particularly when multiple related scoring dimensions are requested simultaneously. In the present study, this behaviour does not affect the final evaluation results, because the two scores contribute equally (1:1 weighting) to the aggregated metric; swapping their positions leaves the combined score unchanged. However, such inconsistencies would become consequential under non-uniform weighting schemes or alternative aggregation strategies.
Table 10: Prompt used for LLM-based scoring of definition correctness and discriminativeness.
Section Prompt Content
Instruction You are a biomedical ontology editor. Score four candidate
definitions for Concept A ONLY on:
- discriminative (1-5): helps distinguish A from concept B.
- correctness (1-5): medically plausible and not misleading.
Output format Return ONLY this JSON:
"scores": [
{{"discriminative": int, "correctness": int}},
{{"discriminative": int, "correctness": int}},
{{"discriminative": int, "correctness": int}},
{{"discriminative": int, "correctness": int}}
Concept A ## Concept A
label: "{a_label}"
definitions:
1) "{d1}"
2) "{d2}"
3) "{d3}"
4) "{d4}"
Concept B ## Concept B
label: "{b_label}"